x=4y+45 Solve the system using the substitution methon
2y+x=-9

Answer by rfer(16322)

2y+4y+45=-9
6y=-54
y=-9
-------------------------------
x=4*-9+45
x=-36+45
x=9
------------------
2*-9+9=-9
-18=-18